# Ferret: Federated Full-Parameter Tuning at Scale for Large Language Models

### v1

@@ C.1 Experimental Setup
Baselines. In line with the comparison in [12], we selected four practical methods for federated LLM tuning as our baselines: (1) FedPTuning [9], (2) FedPrompt [29], (3) FedIT [28], and (4) FedIT-SGD, a variant of FedIT that replaces Adam with SGD. In addition, we included four full-parameter tuning methods for comparison: (1) FedAvg [15], (2) FedZO [20], (3) FedMeZO, a hybrid of FedAvg and MeZO [41], and (4) FedKSeed [12].
@@
Datasets. We conducted our experiments using the Natural Instructions (NI) [26] and Dolly-15K [27] datasets, following a setup similar to [12]. For the NI dataset, we allocated 738 training tasks to individual clients for local updates and reserved 119 test tasks for global evaluation, reflecting a non-IID distribution. Meanwhile, for the Dolly-15K dataset, the final task was utilized for global evaluation, while the remaining tasks were distributed among 200 clients with varying levels of label distribution skew. Rouge-L [42] was chosen as the evaluation metric. Given our resource constraints, we selected DataJuicer-1.3B [24] and LLaMA-3B [25] as the base models for our study.
FL Setting. In each round of federated learning, 5% of clients were randomly selected to participate. Following the same practice in FedKSeed [12], we set the total number of communication rounds to 40 for the NI dataset and 60 for Dolly-15K for all baselines. Due to the compelling efficiency of our method, we set the total number of communication rounds to 12 for the NI dataset and 20 for Dolly-15K for Ferret. First-order baselines trained locally for one epoch, and FedKSeed trained for 200 steps, while our Ferret algorithm trained for 10 iterations (i.e., $T=10$ in Algo. 1). The $K$ value was set to 4096 for FedKSeed. All approaches perform local update with a batchsize of 1 to reduce memory consumption. For each local update iteration in Ferret, we accumulate the gradients from 4 samples.
@@
Hyper-parameters. For Ferret, the local update learning rate $\eta$ for each client is set to $1\times 10^{-4}$, where the selected learning rate is searched from $[2\times 10^{-4},1\times 10^{-4},5\times 10^{-5}]$. The global aggregation learning rates on Natural Instruction and Dolly-15K are set to $10.0$ and $3.0$, respectively, which is search from $[10.0,5.0,1.0]$. For other baselines in Tab. 1 of our main paper, we reported their accuracy performances using the results from FedKSeed [12].
C.2 Calculation of Computational Cost and Communication Overhead
Computation. For FedZO, we use the default hyper-parameters ($b_{1}=20,b_{2}=25$) from the FedZO paper [20], which employs 20 local update steps and 25 perturbations for each sinlge update step. For calculating the computational cost of FedAvg and Ferret, we apply 10 local update steps for each client. Same as our experimental setting in Tab. 1, the batch is set to 1 for all methods. The time cost incurred at gradient projection is also included in the Local Update.
For the overall computation cost, we follow the calculation below:
[…]
Communication. The per-round communication cost refers to the total number of parameters exchanged between a client and the central server during a single round. This includes both the raw or projected gradients that the client sends to the server and the aggregated gradients that the client receives from the server. Each parameter (or projected gradient) is encoded as 16-bit floating point numbers. In accordance with the practice in FedKSeed, we set the number of rounds $R$ to 40 for both FedZO and FedKSeed. Given the notable convergence rate of Ferret, we set $R$ to 12 for both Ferret and FedAvg. Although [12] employs $R=40$ for FedAvg, we use $R=12$ to provide a strong basis for comparison and to highlight the computational efficiency of Ferret.
For the overall communication cost, we follow the calculation below:
[…]
C.3 Prompt Template
@@
In our experiments, the raw input data is pre-processed to follow a structured format, where we warp the input text to the Alpaca prompt template [43]. The corresponding templates for the NI and Dolly-15K dataset are shown in Table 4 and 5.
@@
Below is an instruction that describes a task, paired with an input that provides further context. Write a response that appropriately completes the request. ### Instruction: {Definition} ### Input: {input} ### Response:
Table 5: Prompt template for Dolly-15K. If some data instances do not have the context
attribute, we will discard the line “### Input: ” in the template.
@@
C.4 Ablation Studies on Reconstruction
@@
C.5 Ablation Studies on Convergence and Generalization
@@
Figure 5: Convergence and generalization of Ferret under varying $K$ on Natural Instructions with DataJuicer-1.3B where $2K_{0}$ corresponds to the communication cost of 7.8$\times$103 per round in Tab. 3.
In Fig. 5, we present the convergence and generalization of Ferret under varying $K$ on the Natural Instructions dataset with DataJuicer-1.3B, using the same experimental setup as described in Appx. C.1. Notably, Fig. 5 shows that: (a) a larger number of random bases (i.e., a larger $K_{0}$) generally leads to improved convergence, while the generalization performance remains comparable; (b) $2K_{0}$ already provides compelling convergence and generalization performance, and further increasing $K$ yields only marginal improvements in convergence; and (c) a slight decrease in generalization performance as $K$ increases is likely due to the reduced regularization effect from noisy gradients.
Refer to caption      Refer to caption
(a) Convergence (b) Generalization & Time Cost
Figure 6: Convergence, generalization, and projection time cost per round of Ferret under varying $L$ on Natural Instructions with DataJuicer-1.3B where $L=194$ is applied in our Tab. 2.
In Fig. 6, we present the convergence, generalization, and projection time cost of Ferret under varying block sizes $L$ on the Natural Instructions dataset with DataJuicer-1.3B, using the same experimental setup as described in Appx. C.1. Notably, Fig. 6 shows that increasing the number of blocks (i.e., a larger $L$) leads to improved convergence and reduced time cost for projection and reconstruction, while the generalization performance remains comparable. This improved convergence is likely due to the logarithmic term in the reconstruction error of our (7), as a larger number of blocks reduces the dimensionality of each block, thereby minimizing reconstruction error. In addition, the reduced time cost aligns with our analysis in Sec. 4.1 and the empirical results shown in Fig. 4(c), further highlighting the efficacy of our block-wise reconstruction method (8).

### v3

@@ C.1 Experimental Setup
In line with the comparison in (Qin et al., 2024), we selected four practical methods for federated LLM tuning as our baselines: (1) FedPTuning (Kuang et al., 2024), (2) FedPrompt (Kuang et al., 2023), (3) FedIT (Zhang et al., 2024a), and (4) FedIT-SGD, a variant of FedIT that replaces Adam with SGD. In addition, we included four full-parameter tuning methods for comparison: (1) FedAvg (McMahan et al., 2017), (2) FedZO (Fang et al., 2022), (3) FedMeZO, a hybrid of FedAvg and MeZO (Malladi et al., 2023), and (4) FedKSeed (Qin et al., 2024).
C.1.1 Setup on the Natural Instruction and Dolly-15K Datasets
Datasets.
@@ C.1.1 Setup on the Natural Instruction and Dolly-15K Datasets
We conducted our experiments using the Natural Instructions (NI) (Wang et al., 2022) and Dolly-15K (Conover et al., 2023) datasets, following a setup similar to (Qin et al., 2024). For the NI dataset, we allocated 738 training tasks to individual clients for local updates and reserved 119 test tasks for global evaluation, reflecting a non-IID distribution. Meanwhile, for the Dolly-15K dataset, the final task was utilized for global evaluation, while the remaining tasks were distributed among 200 clients with varying levels of label distribution skew. Rouge-L (Lin, 2004) was chosen as the evaluation metric. Given our resource constraints, we selected DataJuicer-1.3B (Chen et al., 2023) and LLaMA-3B (Touvron et al., 2023a) as the base models for our study. The corresponding HuggingFace model paths are “datajuicer/LLaMA-1B-dj-refine-150B” and “openlm-research/open_llama_3b”.
Hyper-parameters.
@@
For Ferret, the local update learning rate $\eta$ for each client is set to $1\times 10^{-4}$, where the selected learning rate is searched from $[2\times 10^{-4},1\times 10^{-4},5\times 10^{-5}]$. The global aggregation learning rates on Natural Instruction and Dolly-15K are set to $10.0$ and $3.0$, respectively, which is search from $[10.0,5.0,1.0]$. For other baselines in Tab. 1 of our main paper, we reported their accuracy performances using the results from FedKSeed (Qin et al., 2024).
Prompt Template.
@@
In our experiments, the raw input data is pre-processed to follow a structured format, where we warp the input text to the Alpaca prompt template (Taori et al., 2023). The corresponding templates for the NI and Dolly-15K dataset are shown in Tab. 7 and 8.
@@
### Instruction: {Definition}
### Input: {input}
### Response:
Table 8: Prompt template for Dolly-15K. If some data instances do not have the context attribute, we will discard the line “### Input: ” in the template.
@@
C.1.2 Setup on the CodeAlpaca and GSM8K Datasets
Datasets.
To further demonstrate that Ferret can also improve the capability of larger LLMs for code generation and mathematical reasoning, we conducted more experiments using the CodeAlpaca (Chaudhary, 2023) and GSM8K (Cobbe et al., 2021) datasets, following a similar federated setup. The CodeAlpaca dataset (of around 8.0k samples) is a code dataset that consists of ten programming languages, including C, C#, C++, Go, Java, PHP, Pascal, Python, Scale, and X86-64 Assemble. We exclude the X86-64 Assembly data due to limited samples in the dataset. We uniformly randomly sampled 10% instances from the original data as the hold-out test set for evaluation, and we split the remaining 10% samples into nine subsets based on the programming language category and assign each subset to one client as its local training data. For GSM8K, its official train set is split into three subsets, where each client’s dataset consists of grade school math questions randomly partitioned from the original dataset, forming a IID distribution. We use the official GSM8K test split as the evaluation dataset. Rouge-L (Lin, 2004) was chosen as the evaluation metric. To demonstrate the scalability of Ferret, we extended the experiments to larger models: LLaMA2-7B and LLaMA2-13B (Touvron et al., 2023a) as the base models for our study. The corresponding HuggingFace model paths are “meta-llama/Llama-2-7b-hf” and “meta-llama/Llama-2-13b-hf”.
Hyperparameters.
For FedZO and FedKSeed, the local update learning rate is set to $3\times 10^{-7}$ for all models. For FedAvg on both LLaMA2-7B and LLaMA2-13B, the local update learning rate $\eta$ for each client is set to $3\times 10^{-4}$, and the global aggregation learning rate is set to $1.0$. For Ferret on LLaMA2-7B, the local update learning rate $\eta$ is set to $3\times 10^{-4}$ and the global aggregation learning rate is set to $5.0$. For Ferret on LLaMA2-13B, the local update learning rate $\eta$ is set to $5\times 10^{-4}$ and the global aggregation learning rate is set to $10.0$. The selected learning rate is searched from $[5\times 10^{-4},3\times 10^{-4},1\times 10^{-4}]$ and the selected global aggregation learning rates is searched from $[10.0,5.0,1.0]$.
C.2 Calculation of Computational Cost and Communication Overhead
Calculation of Computational Cost.
For FedZO, we follow the same hyper-parameters ($b_{1}=200,b_{2}=1$) for FedZO from FedKSeed paper (Qin et al., 2024), which employs 200 local update steps and 1 perturbation for each local update step. For calculating the computational cost of FedAvg and Ferret, we apply 10 local update steps for each client. Same as our experimental setting in Tab. 1, the batch size is set to 1 for all methods. The time cost incurred at gradient projection is also included in the Local Update.
For the overall computation cost per round, we follow the calculation below:
[…]
Calculation of Communication Overhead.
The per-round communication cost refers to the total number of parameters exchanged between a client and the central server during a single round. This includes both the raw or projected gradients that the client sends to the server and the aggregated gradients that the client receives from the server. Each parameter (or projected gradient) is encoded as 16-bit floating point numbers. In accordance with the practice in FedKSeed, we set the number of rounds $R$ to 40 for both FedZO and FedKSeed. Given the notable convergence rate of Ferret, we set $R$ to 12 for both Ferret and FedAvg. Although (Qin et al., 2024) employs $R=40$ for FedAvg, we use $R=12$ to provide a strong basis for comparison and to highlight the computational efficiency of Ferret.
C.3 More Comparison on LLaMA3-8B and Qwen2.5-7B Models
We conducted additional experiments on CodeAlpaca and GSM8K using LLaMA3-8B and Qwen2.5-7B in Tab. 9. The results demonstrate the consistent effectiveness of our Ferret, achieving near FedAvg performance with significantly reduced communication overhead across models and tasks.
Table 9: More comparison of Rouge-L (%) among various algorithms on LLaMA3-8B and Qwen2.5-7B models. Each cell reports the mean $\pm$ std of Rouge-L scores from the final round of four runs, each with a different random seed. The highest and second-highest scores are shown in bold and underline, respectively.
Algorithm CodeAlpaca GSM8K
LLaMA3-8B Qwen2.5-7B LLaMA3-8B Qwen2.5-7B
FedZO 16.66 $\pm$ 0.50 9.14 $\pm$ 0.32 7.79 $\pm$ 1.36 23.84 $\pm$ 1.19
FedKSeed 5.73 $\pm$ 1.26 9.14 $\pm$ 0.32 7.79 $\pm$ 1.36 23.84 $\pm$ 1.19
FedAvg 19.88 $\pm$ 0.67 17.47 $\pm$ 0.49 45.48 $\pm$ 0.51 43.86 $\pm$ 0.36
Ferret (ours) 19.59 $\pm$ 0.66 14.64 $\pm$ 0.74 45.07 $\pm$ 0.78 38.28 $\pm$ 1.70
C.4 More Comparison of Computational Cost and Communication Overhead
[…]
Table 10: Comparison of computational cost and communication overhead on LLaMA2-13B, focusing on (a) the computational costs from local updates, global aggregation, and the overall tuning process; and (b) the per-round and overall communication costs. The improvement achieved by our Ferret is reported in brackets using blue (compared with FedKSeed) and orange (compared with FedAvg).
Algorithm Computational Cost (Sec.) Communication Cost
Local Update Global Aggr. Overall (# params.)
FedZO 114.1 25.7 139.8 2.6$\times$1010
FedKSeed 188.4 666.2 854.6 8.2$\times$103
FedAvg 24.9 25.7 50.6 2.6$\times$1010
[…]
C.5 Ablation Studies on Reconstruction
@@
Comparison of Reconstruction Accuracy between Ferret and ZO Method.
@@
C.6 More Ablation Studies on Convergence and Generalization
@@
Figure 9: Convergence and generalization of Ferret under varying optimizers for local updates.
Convergence and Generalization of Ferret under Varying Optimizers.
In Fig. 9, we present the convergence and generalization of Ferret under different optimizers for its local updates, using the same experimental setup described in Appx. C.1. Notably, Fig. 9 demonstrates that Ferret achieves faster convergence with an improved optimizer (e.g., Adam vs. SGD) while maintaining comparable generalization performance. These findings further support the adaptability of Ferret, as discussed in Sec. 4.3.